The Labrador Falls Company has three divisions: A Division, B Division, and C Division.

A B C
Sales $ 323,000 $ 543,000 ?
Net operating income 61,500 ? $ 24,300
Residual income ? 37,500 14,700
Average Division Assets ? ? 81,500
Cost of Capital 12 % 10 % ?
Profit Margin 20 % 4 % ?
Asset Turnover ? 1 ?
Return on investment 15 % ? ?


What was A Division's residual income last year?
Answer #1

Residual income = Net operating income - Minimum operating income

Average operating assets = 61500*100/15 = 410000

Residual income A = 61500-(410000*12%) = 12300